OpenStudy (anonymous):

Find the dot product, a • b. a = 6i + 5j, b = -5i + 4j

OpenStudy (anonymous):

multiple the i with the i and the j with the j since i . i =1 n j . j =1 you will get the answer in magnitude form...

OpenStudy (anonymous):

-30,20 ?

OpenStudy (anonymous):

\[a \cdot~b=(6i+5j)~\cdot~(-5i+4j)\]

OpenStudy (anonymous):

correct so u have \(a \cdot~b=-30+20=?\)

OpenStudy (anonymous):

-10!

OpenStudy (anonymous):

Perfect :) All THe best :)